OpenStudy (anonymous):

three and two-thirds plus 4 and one-eighth equals 7 and three-elevenths what did i do wrong

OpenStudy (anonymous):

what

OpenStudy (mathstudent55):

\( 3\dfrac{2}{3} + 4\dfrac{1}{8} = 7\dfrac{3}{11} \)

OpenStudy (anonymous):

Yea, I was typing that.

OpenStudy (anonymous):

what is that

OpenStudy (mathstudent55):

When you add fractions, you need a common denominator. Then you add the numerators and keep the same common denominator. Here, the fractions were "added" by adding the numerators and adding the denominators, which is incorrect.

OpenStudy (mathstudent55):